# CIF-T: A Novel CIF-based Transducer Architecture for Automatic Speech Recognition

### v3

@@ 4.2 Results
Table 2 shows the character error rate (CER) results on the AISHELL-1 dataset. We use CTC loss and LM loss on the vanilla RNN-T model as our baseline model ${}^{\ddagger}$. When just employing the CIF mechanism to replace the alignment process in the ${}^{\ddagger}$, ${}^{\ddagger}$ achieves a competitive result of 4.6%/5.3% on the dev/test set compared to 4.7%/5.3% for the ${}^{\ddagger}$ as a baseline. With the improvements proposed in this work, the CIF-T(S) and CIF-T(M) achieve consistent or superior performance compared to the best publicly available Conformer-based AED models, CIF-based models, and RNN-T-based models, with equal or smaller parameters. Furthermore, the CIF-T(L) achieves a state-of-the-art result of 4.1%/4.3% on the dev/test set. We also evaluate the results of applying UGBP on ${}^{\ddagger}$, and despite the better fusion method improved performance, the results still fall behind our CIF-T models. The CER results on the WenetSpeech are presented in Table 3. CIF-T achieves a significantly lower CER results than ESPnet and Wenet with essentially the same parameters, while achieving a competitive results compared to Conformer-MoE which has more than three times parameters.

### v4

@@ 4.2 Results
Table 2 shows the character error rate (CER) results on the AISHELL-1 dataset. We use CTC loss and LM loss on the vanilla RNN-T model as our baseline model RNN-T. When just employing the CIF mechanism to replace the alignment process in the RNN-T, CIF-T achieves a competitive result of 4.6%/5.3% on the dev/test set compared to 4.7%/5.3% for the RNN-T as a baseline. With the improvements proposed in this work, the CIF-T(S) and CIF-T(M) achieve consistent or superior performance compared to the best publicly available Conformer-based AED models, CIF-based models, and RNN-T-based models, with equal or smaller parameters. Furthermore, the CIF-T(L) achieves a state-of-the-art result of 4.1%/4.3% on the dev/test set. We also evaluate the results of applying UGBP on RNN-T, and despite the better fusion method improved performance, the results still fall behind our CIF-T models. The CER results on the WenetSpeech are presented in Table 3. CIF-T achieves a significantly lower CER results than ESPnet and Wenet with essentially the same parameters, while achieving a competitive results compared to Conformer-MoE which has more than three times parameters.